OpenStudy (anonymous):

f(x)=(1/9)x^9 -a^(8) x. Find the critical point, assuming a is a constant.

OpenStudy (anonymous):

The a variable messes me up, not sure how to approach this. Screenshot attached.

OpenStudy (campbell_st):

whats the 1st derivative... and whats the 2nd derivative?

OpenStudy (anonymous):

err how do i do that with a though?

OpenStudy (anonymous):

The first part is x^8, but the next part not sure. Do i need product rule or something?

OpenStudy (campbell_st):

ok... so \[f'(x) = x^8 - a^8\] \[f"(x) = 8x^7\] set the 1st derivative to zero and solve \[0 = x^8 - a^8\] are you able to solve the derivative..?

OpenStudy (anonymous):

Oh i was over complicated it. That makes sense.

OpenStudy (anonymous):

Do i solve for x or a though?

OpenStudy (anonymous):

\[x=\sqrt[8]{a^8} \] Is this correct?

OpenStudy (campbell_st):

so you have 2 stationary points from the 1st derivative... use the 2nd derivative to test the nature... and solve the 2nd derivative.... for a point of inflection... then check for a change in concavity..

OpenStudy (campbell_st):

well you are on the way... what values can x take...?

OpenStudy (campbell_st):

if \[x = \sqrt[8]{a^8}...then... x = \pm a\]

OpenStudy (campbell_st):

I'll leave you to finish... good luck

OpenStudy (anonymous):

Right so is that the answer, i don't see what else we could do? x=a?

OpenStudy (anonymous):

Seems pointless that x just =a.

OpenStudy (anonymous):

or +a, -a

OpenStudy (campbell_st):

ok... so you have the critical points or stationary points.. there is also 1 in the 2nd derivative \[0 = 8x^7\] but given you don't need to determine there nature... from my point of view the critical points are x = -a, x = 0 and x = a the values of -a and a are max or mins... x = 0 is a possible point of inflection.

OpenStudy (anonymous):

Hmm okay that makes sense i guess.